OpenStudy (anonymous):

HELP! will fan and medal. Solve 5

OpenStudy (hauntedwoodsgal):

EXAMPLE 1. Given f(x) = 2x + 3. Find the values of a) f(0), b) f(7), c) f(-5). Solution: a) f(0) means that x = 0. Replace the x with the value of 0. f(x) = 2x + 3 f(0) = 2(0) + 3 = 3 b) f(7) means that x = 7. Replace the x with the value of 7. f(x) = 2x + 3 f(7) = 2(7) + 3 = 14 + 3 or 17 c) f(-5) means that x = -5. Replace the x with the value of -5. f(x) = 2x + 3 f(-5) = 2(-5) + 3 = -10 + 3 or -7